### v1

Abstract
Machine learning (ML) is a rapidly developing area of medicine that uses significant resources to apply computer science and statistics to medical issues. ML’s proponents laud its capacity to handle vast, complicated, and erratic medical data. It’s common knowledge that attackers might cause misclassification by deliberately creating inputs for machine learning classifiers. Research on adversarial examples has been extensively conducted in the field of computer vision applications. Healthcare systems are thought to be highly difficult because of the security and life-or-death considerations they include, and performance accuracy is very important. Recent arguments have suggested that adversarial attacks could be made against medical image analysis (MedIA) technologies because of the accompanying technology infrastructure and powerful financial incentives. Since the diagnosis will be the basis for important decisions, it is essential to assess how strong medical DNN tasks are against adversarial attacks. Simple adversarial attacks have been taken into account in several earlier studies. However, DNNs are susceptible to more risky and realistic attacks. The present paper covers recent proposed adversarial attack strategies against DNNs for medical imaging as well as countermeasures. In this study, we review current techniques for adversarial imaging attacks, detections. It also encompasses various facets of these techniques and offers suggestions for the robustness of neural networks to be improved in the future.
keywords:
Adversarial attack, Medical image, Deep Neural Network, Model safety, Robustness
1 Introduction
1.1 Adversarial Attack
Adversarial attacks encompass a set of techniques aimed at manipulating machine learning models by introducing well-crafted, often imperceptible alterations to input data. These modifications aim to deceive the model, leading to misclassifications or inaccurate predictions [1]. Adversarial attacks have the potential to compromise machine learning systems’ dependability and security, which is concerning for crucial applications such as cybersecurity, autonomous vehicles, and medical diagnosis. Of all the adversarial attack types, ”White-box attacks” and ”Black-box attacks” are the most common.
Ongoing research in machine learning security centers on the analysis of adversarial attacks and the development of defense strategies. Researchers continually explore novel attack tactics to identify potential weaknesses in AI systems while concurrently striving to construct robust models less susceptible to adversarial perturbations. Artificial intelligence (AI) modern technology is widely successful in fields like computer vision, natural language processing, and automated driving; however, its application in vital safety sectors is hindered by its susceptibility to adversarial attacks. Consequently, enhancing the resilience of AI systems against such attacks has become paramount for the progress of AI [2].
The rapid advancement of AI technologies has found diverse applications in numerous domains, from machine translation, speech recognition, and object identification to more intricate tasks like drug composition analysis [3, 4, 5, 6, 7]. Noteworthy applications include brain circuit construction [8], particle accelerator data analysis [9], [10], and DNA mutation impact analysis [11]. Since Szegedy et al.’s seminal work [12] highlighting neural networks’ vulnerability to adversarial attacks, research on adversarial technologies for artificial intelligence has steadily expanded. New techniques for adversarial attacks and strategies for mitigating them are consistently emerging.
The term ”adversarial attacks at the training stage” refers to actions taken by adversaries to manipulate the training dataset, input characteristics, or data labels during the training phase of the target model. This manipulation of the training dataset involves actions such as adding or deleting training data, as demonstrated by Barreno et al.’s approach [13]. During the testing phase, adversarial assaults are categorized into white-box attacks and black-box attacks [2]. In white-box situations, attackers have access to the parameters, techniques, and structure of the target model, enabling them to craft adversarial samples based on this information.
1.2 The Adversary’s Objective: Evasion attack versus poisoning assault
”Poisoning attacks” are assault methods that allow an attacker to add or change a large number of fake samples to the training set of a DNN algorithm. The trained classifier may perform poorly as a result of these bogus data. They may have poor accuracy [14] conversely make imprecise predictions on a few samples for analysis [15]. The classifiers used in evasion attacks are fixed and often work well on safe testing samples. The adversaries are unable to reform the classifier’s settings or parameters, but they do generate some fictitious instances that the classifier is unable to distinguish.
1.3 Deep Neural Networks: Adversarial Attacks
Regarding medical image processing use cases such as diagnosing cancer and lesion identification, deep neural networks, also known as DNNs, are rapidly gaining popularity [16]. Nonetheless, a new study suggests that adversarial examples/attacks with tiny, barely noticeable changes can weaken medical deep learning systems. There are presently safety concerns associated with the usage of these devices in clinical settings [16]. Deep neural networks (DNN) are increasingly well-liked and effective at a variety of machine learning requisitions. They have been employed with surprising effectiveness in a variety of recognition issues in the fields of pictures, graphs, text, and voice [1]. They are able to identify things in images with accuracy that is almost human [17], [18]. Additionally, they are employed in speech recognition [19], natural language processing [20], and gaming [21].
Figure 1: The attack of Biggio’s SVM classifier for letter recognition. [1]
Regarding the data set provided by MNIST, Biggio et al.[22] produce adversarial instances first, focusing on traditional machine learning classifiers like SVMs and 3-layer neural networks with full connectivity which is shown in Figure 1. In order to trick the classifier, it optimizes the discriminant function. As an illustration, consider a linear SVM classifier on the MNIST dataset.
1.4 Examining examples of opposition in the real world
The research [23] put decals to traffic signs that pose a serious threat to autonomous vehicles’ sign recognition technology. These hostile objects are especially detrimental to deep learning systems since they can interfere directly with numerous real-world DNN usages, such facial recognition and autonomous cars. By assessing the adversarial images (FGSM, BIM) produced to see if they are ”robust” against changes in natural circumstances (e.g., shifting viewpoint, illumination, etc.), the authors of the work [24] investigate the viability of creating tangible adversarial objects. Robust in this context means that even after transformation, the produced pictures are still antagonistic. The experiment’s findings show that many of these adversarial examples—particularly those produced by FGSM—remain antagonistic to the classifier even after transformation. The findings imply that antagonistic real-world objects could trick the sensor in many situations.
1.5 Medical Image Under Artificial Intelligence Attack
Researchers have access to strong models of developing science and technology thanks to deep learning. Convolutional neural networks, also known as CNNs, are among the most significant categories of deep learning (DL) frameworks for the analysis and processing of images because of their exceptional ability to learn valuable information. The processing of the human organism utilizing different picture modalities for therapeutic, diagnostic, and health surveillance purposes is known as medical image analysis [25]. The advancement of computer vision through the use of deep neural networks addresses issues that were not well-solved by traditional image processing methods. Beinfeld et al. [26] asserted that \$385 spent on medical imaging results in a savings of almost \$3000. MRI, CT scans, ultrasound (US), and X-rays are the most frequently used image modalities. Figure 2 shows how adversarial assaults may be used to arbitrary alter diagnosis outcomes in three different medical picture datasets: Dermoscopy [27], X-ray of the chest [28], and Fundoscopy [29].
Figure 2: Instances of adversarial approaches designed by the PGD (Projected Gradient Descent) to deceive DNNs developed using datasets of medical images include dermoscopy [27] (third row), chest x-ray [28], and fundoscopy [29] (first row, DR=diabetic retinopathy). Normal images on the left, adversarial perturbations in the middle, and adversarial images on the right. The left-bottom tag indicates the projected class, and green or red indicates whether the predictions were accurate or not [16].
Hu et al. [1] assert that in order to develop more reliable models, it is crucial to investigate the reasons why adversarial cases emerge and to comprehend deep learning models better. Depending on the knowledge of the enemy, attacks can be categorized into three kinds. The saliency (or attention) map of a picture that was input shows the regions that significantly change the model’s output based on the gradients of the categorization loss with regard to the input [16]. We can see that some medical photos have highly concentrated regions that are noticeably larger. This could indicate that the rich biological textures included in medical photos occasionally entice the DNN model to focus more on aspects unrelated to the diagnosis. Tiny adjustments in these high-focus areas can significantly affect the model’s output.
Due to several factors, including High-Dimensional Data, High-Dimensional Data, Transferability of Adversarial Examples, Complex Decision Boundaries, Black-Box Attacks, Safety-Critical Applications, and Safety-Critical Applications, Medical Image Deep Neural Network (DNN) models can be relatively easier to attack when compared to some other domains. Despite these difficulties, experts in the field are actively attempting to create more reliable and safe DNN models for medical images. To make these models more resistant to hostile attacks, strategies like adversarial training, input preprocessing, and defensive distillation are being investigated. To further enhance the security and dependability of AI systems in healthcare applications, consistent evaluation frameworks for adversarial robustness in medical imaging must be developed.
Byra et al. [30] proposed an attack strategy on ultrasonography (US) imaging for liver fatty tissue. Radio-frequency signals are used to rebuild US pictures, and the reconstruction technique was subjected to a zeroth-order optimization attack [31]. The InceptionResNetV2 model was used in the studies, and the assault resulted in a 48% loss in model accuracy. Ozbulak et al. [32] proposed an attack specifically designed for medical picture segmentation called the adaptive segmentation masking attack (ASMA). The suggested attack offers significant intersection-over-union (IoU) degradation and produces nearly undetectable samples for most portions. Because the U-Net framework is among the most well-known models for clinical picture segmentation, they employed it in the trials. The datasets used were for segmentation of glaucoma optic disk [33] and ISIC skin lesion segmentation [34]. Chen et al. published a method for fabricating hostile cases to thwart medical picture segmentation [35]. In order to simulate anatomical and intensity fluctuations, geometrical deformations are used to create the adversarial examples. By attempting to partition organs from abdominal CT scans using a U-Net model, they tested the effectiveness of these examples. With reference to the Dice score measure, they successfully reduced it significantly across all organs. The kidneys and pancreas, however, require a higher level of disturbance and are more arduous to assault than the liver as well as spleen.
Figure 3: Clean image, bias field noise, and diagnosis following application of bias field noise are shown in (a), (b), and (c), respectively [36].
Tian et al. [36] examined the phenomena of bias field, which can arise from incorrectly capturing a medical image and jeopardize a DNN’s efficacy, as shown in Figure 3. Motivated by adversarial attacks, the authors created an adversarial-smooth bias discipline approach to trick a model. The ResNet50, MobileNet, among others and DenseNet121 models were employed to fine-tune the chest X-ray dataset that was used in their study. They looked at this attack’s transferability and white-box attacks. Comparing the suggested attack to other cutting-edge white-box attacks, it exhibited a greater attack accuracy on transferability.
This review’s major organization is as follows: We offer some significant adversarial notions on medical images in Section 2. Additionally, it provides a detailed overview of earlier research. We explore a few strategies in Section 3 with regard to the picture classification scenario. We utilize Section 4 to quickly summarize some findings from earlier research that tries to explain the phenomena of antagonistic examples on medical data. The overview is concluded in Section 5.
2 Background Literature Review
Adversarial instances are intentionally created data inputs that are intended to degrade a machine learning model’s performance. When researchers examined strategies used by spammers to evade spam filters in 2004, they unofficially coined the term ”adversarial inputs” [37]. These misleading examples are usually produced by deliberately faking real data, such spam advertising messages, in order to trick the computer system that analyzes it. Alterations can be applied to text data, like spam, by introducing innocuous text or modifying frequently used phrases in malicious communications with synonyms. To fortify algorithms against adversarial attacks, researchers have explored various strategies, including training algorithms on adversarial samples and employing sophisticated data processing techniques to minimize the susceptibility to manipulation. The quest for fully robust models in machine learning aims to accelerate the development of algorithms capable of making decisions based on consistent explanations, with promising early efforts in this direction [38].
In the realm of healthcare, medical claims codes play a crucial role in determining the amount spent on a patient visit post-payer approval. Payers often assess these claims using automated fraud detectors increasingly driven by machine learning. Historically, healthcare providers have shaped payer records of patient visits, including the associated codes, to influence the algorithmic outputs of payers [38]. Medical fraud, a market valued at \$250 billion, exemplifies the extreme end of this strategic tailoring of patient presentations. While some practitioners may overtly fabricate medical claims, patient data falsification often takes more covert forms. For instance, consistently submitting codes for billing services identical to, but more expensive than, those actually provided is known as intentional upcoding. Here, Table 1 shows past exclusive past research works where authors did some fantastic work for medical images to defend adversarial attacks. Different researcher proposed different frameworks to overcome this problem and to suggest a new thought for revolution.
2.1 Elements of a Cyber Adversary’s Assault
Figure 4: A showcase of executing adversarial attacks on various medical AI systems without resorting to overtly deceptive data manipulation [39].
In some instances, subtle adjustments to billing codes blur the boundary between fraudulent activities and conscientious best practices. One notable illustration may be found in the guidelines on the Endocrine Society website, which advise medical professionals not to bill for metabolic syndrome, or International Classification of Diseases (ICD) code 277.77, in situations of obesity. Coverage denial is likely to occur with this particular code/condition combo in Figure 4. The Society recommends an alternative approach, suggesting billing for the individual codes related to specific disorders comprising the metabolic syndrome, such as hypertension.
Table 1: Existing adversarial attack research on medical Image
Reference Year Framework Contribution
[40] 2021 Propose a framework to defend against adversarial training, speckle-noise attacks, and maintain accurate classification labels in recognizing diabetic retinopathy through analyzing retinal fundus images. a) accuracy of 99%; b) defensive model’s robustness ; c) a system that includes a fresh SN attack.
[41] 2022 Explain the use of gradient-free trained sign activation systems in medical imaging AI systems to identify and thwart hostile attacks. The model performs better and even twice as well as the finest.  a) Owing to significant distortion and excellent transferability accuracy, the model’s average success rate in adversarial case classification is 88.89%; b) This model defends against adversarial attacks.
[42] 2020 This method detects diverse adversarial attacks while preserving user anonymity and classification performance, making it a versatile addition for deep learning-based medical imaging systems to enhance resilience. a) The chest X-ray dataset exhibits robust performance across diverse configurations, contributing to the heightened safety of deep learning-based medical image classification systems; b) To improve medical image categorization systems based on deep learning security.
[43] 2022 In weakly-supervised clinical tasks, CNNs and ViTs exhibit vulnerability to both white- and black-box adversarial attacks with comparable baseline performance. a)Compared to CNNs, aViTs have a larger latent portrayal of clinically meaningful categories; b) Findings align with prior theoretical investigations and provide tangible evidence of ViTs’ capacity to grasp computational pathology; c) This suggests that computational pathology AI models will be deployed widely.
[…]
2.2 The components of an adversarial attack
Figure 5: Distribution of chosen articles on a yearly basis Pie chart; line graph [44].
Presenting a thorough overview of the various adversarial attack strategies and defense techniques is the aim of the study [44]. We first discuss the theoretical underpinnings, practices, and practical uses of adversarial attack strategies. Then, a brief discussion of the research on defense strategies obscuring the boundary of the large field follows. A few selected articles are published year-by-year from 2012 to 2021, as seen in Figure 5. This work aims to provide thorough taxonomies, evaluations of harmful assaults, and protections against the full DL pipeline. In this context, the numerous attack and defense tactics that have been developed over the previous two years have been categorized, with a focus on the clinical deep learning systems that are susceptible to adversarial attacks.
Several defensive strategies have been put forth in an effort to neutralize possible threats. A popular method in natural imaging called adversarial training adds adversarial images to the training dataset in order to make Convolutional neural network, or CNN, models more resilient. Nevertheless, this method is not the best for datasets pertaining to medical imaging, since the addition of different adversarial images could considerably reduce classification accuracy. A robust detection technique for malicious images is presented in the work of Li et al. [42], successfully defending against attacks on deep learning-driven medical image labeling systems. In a different study [43], scientists compared CNN performance with Vision Transformers (ViTs) to determine how resilient CNNs are to different types of attacks in computational pathology. The authors developed robust neural network models, evaluating their efficacy against both white- and black-box attacks. The structures of attacks for both models were scrutinized, with an exploration of the underlying factors influencing their performance. The study’s findings were validated through two clinically relevant classification tasks involving distinct patient groups. Preceding Ma et al.’s research [16], ambiguity surrounded medical image adversarial attacks (AAs), limiting adversarial machine learning analyses to natural images. Unlike natural images, medical images may contain domain-specific elements, impacting medical deep-learning systems susceptible to AAs. AAs possess the potential to manipulate diagnoses and outcomes, emphasizing the critical need for a robust healthcare infrastructure to mitigate potential risks from malicious attacks.
3 Materials and Methods
3.1 Medical Adversarial Defensive Strategies
Various defense models, including techniques such as input denoising, input gradients regularization, and adversarial training, have been devised. However, recent attacks often manage to evade or partially circumvent these protective measures. This paper undertakes a comprehensive exploration of adversarial attack and defense methods within the domain of medical image analysis. It introduces a novel taxonomy based on the application context for a family of techniques, as outlined in Dong et al.’s work [45]. The author establishes a unified theoretical framework encompassing diverse adversarial attack and defense strategies tailored for medical imaging applications.
The primary emphasis in most research efforts on adversarial attacks in medical image processing is on the white-box scenario. These studies, characterized by a comprehensive understanding of medical Deep Neural Networks (DNNs), concentrate on the susceptibility of computer-aided diagnosis models across various medical imaging applications. In the execution of adversarial attacks, attackers may utilize the target diagnosis DNN as a locally deployed model. In addition to white-box adversarial attacks for medical classification tasks, scholars have delved into vulnerabilities associated with various medical imaging tasks. Notably, these publications focus predominantly on medical segmentation, as evident in works such as Chen et al.’s [35] and Ozbulak et al.’s [32].
For natural photos, semi-white-box (Gray-box) attacks have been extensively researched [46]. However, there are only a few academic publications [47], [48] that address this attack scenario for medical image processing. The semi-white-box adversarial approach typically consists of two stages: 1) To create adversarial instances against the target DNN model, the attacker trains a generative model. The attacker has complete access to the target model during training, including backward propagation gradients. 2) Instead of needing to know anything about the target model, as would be the case in a completely black-box scenario, the adversary generator can directly obtain adversarial examples against the target model with the input of authentic photos during the application stage.
Presently utilized white-box adversarial attacks often require multiple backward gradients of the target model. In simpler terms, attackers generate comparable adversarial samples by treating the target Deep Neural Network (DNN) as if it were a locally deployed model. However, due to its reliance on an in-depth understanding of the DNN model to execute an attack, the white-box approach may be unreliable in real-world scenarios. Conversely, the general black-box scenario may offer a more suitable environment for simulating real-world adversarial attacks, with numerous proposals focusing on investigating black-box assaults for natural images.
The outlined solutions usually involve multiple queries to the black-box model or depend on having a comprehensive understanding of the desired diagnosis model. However, in many real-world situations, attackers may lack direct access to the target diagnosis model. The restricted black-box (no-box) configuration, specifically representing the most challenging scenario for real-world adversarial assaults, may be more covertly dangerous even without querying the target black-box DNN. The transferability [49] of adversarial instances across different DNN models fundamentally determines the success of no-box attacks. For example, a no-box attacker can generate adversarial images based on a locally deployed surrogate model, transferring them directly to target medical diagnosis systems. Restricted black-box adversarial attacks are deemed more subtly perilous for natural vision tasks according to a renowned study.
Considering the substantial risks to the healthcare sector, numerous defense strategies have been proposed to counter medical adversarial attacks. In order to overcome the weaknesses shown by adversarial examples, a number of methods have been developed to provide trustworthy deep learning-based systems for natural pictures [49]. The development of accurate computer-aided diagnosis models for clinical applications not only contributes to trustworthy healthcare services for millions but also underscores the importance of investigating adversarially robust models within the realm of medical image analysis.
In their work [45], the authors succinctly summarize studies on adversarial defense in the context of medical image analysis. Various attack methods in the realm of medical images serve as robustness evaluation criteria for adversarial defense, complementing the extensive range of adversarial attack methods mentioned. The minimal separation between adversarial examples and the model’s decision boundary signifies the model’s resistance to disturbances. The goal of adversarial training is to mitigate subsequent losses.
Minimize the
$L(f(x_{\text{adv}},y)+\lambda\cdot L(x,x_{\text{adv}}))$ (1)
where $L(f(x_{\text{adv}},y)$ is the classification loss, $L(x,x_{\text{adv}})$ is the perturbation size, and $\lambda$ is a trade-off hyperparameter.
Transforming the input data into an alternative feature space that diminishes the impact of adversarial perturbations on the model. A technique aimed at enhancing resilience against adversarial attacks involves introducing noise to the features or intermediate representations that characterize the input data.
3.2 Adversarial Training
Most medical adversarial defense techniques focus on using adversarial training to create reliable diagnoses systems. A significant fraction of the works among them go beyond current techniques for natural image adversarial training to tasks relating to medical classification [50], [51], Vatian et al. [52] looked into opposing examples for medical imaging and tried a number of defense strategies to oppose these nefarious representations.
3.3 Adversarial Detection
Adversary detection seeks to identify adversary cases from input examples during the application stage, as opposed to developing robustness during the training stage of computer-aided diagnosis models.Many adversarial detection techniques have been put forth in the field of medical image analysis to stop additional misdiagnosis caused by adversarial samples [42]. In particular, it is possible to think about medical adversarial detection as an anomaly detection issue that can be resolved by combining explainability approaches [53].
3.4 Image-level Pre-processing
A clean image and the related adversarial perturbation make up an adversarial image in general. Meanwhile, it has been shown that DNNs can perform well on clean images while still being vulnerable to adversarial examples [12]. Denoising the adversarial example to remove the perturbation component can therefore help make the subsequent network diagnostic easier. Image-level pre-processing can be useful and secure in the context of biomedical image analysis because it does not require re-training or modifying medical models.
3.5 Improvement of Features
The difference in robustness between human and machine vision is attributed to adversarial examples linked to non-robust features extracted from specific patterns in the data distribution [54]. Consequently, enhancing feature representation is crucial for the development of robust inference systems. In this investigation, we characterize feature augmentation as the modification of architectures or mapping functions. Various techniques aimed at improving features have significantly enhanced the robustness of medical classification models [55].
3.6 Distillation of Knowledge
In the field of machine learning, knowledge distillation is a helpful technique for transferring learned information from a complex (teacher) model to a simple (student) model. Therefore, the situation where the network structures for the teacher and learners are identical is specifically referred to as self-distillation. Furthermore, a great deal of research has been done on the natural imaging domain using adversarial knowledge distillation, which moves the adversarial resilience from a heavy teacher model to a view student model [56].
4 Result and Discussion
This study utilizes four publicly accessible benchmark datasets, as detailed in [45], to investigate adversarial attack and defense in the context of medical image processing (illustrated in Figure 6). Firstly, the Messidor1 dataset comprises 1,200 eye fundus color images for detecting diabetic retinopathy across four classes based on retinopathy grade. Secondly, the International Skin Imaging Collaboration (ISIC) dataset consists of 2,750 dermoscopic images categorized into three classes for skin lesion classification and segmentation. Thirdly, the ChestX-ray 14 dataset comprises 112,120 frontal-view X-ray images representing 14 thorax diseases. Lastly, the COVID-19 database incorporates 21,165 chest X-ray images accompanied by segmentation-capable lung masks.
Figure 6: Medical adversarial examples with predictions for a range of perturbation sizes. For visualization, the created segmentation masks are placed on top of the source photos [45].
The study conducted by the authors [57] employs experiments to demonstrate the following: 1) The SSAT module significantly enhances the adversarial robustness of the model without compromising the classification accuracy of clean images. 2) The UAD module effectively identifies and excludes a majority of successful adversarial examples. 3) In comparison to other existing AI systems, their medical imaging AI solution (UAD + SSAT) minimizes the risk of adversarial attacks. The research utilizes a publicly available dataset of retinal OCT images, employing the ”white-box” setting as the most challenging threat for assessing class prediction performance. Across all white-box attack conditions, the authors establish that SSAT consistently outperforms alternative baselines while maintaining comparable or superior performance for clean image classification. Furthermore, the authors demonstrate that the combined approach of UAD complementing SSAT results in the lowest adversarial risk according to the new measure presented. Regardless of the training techniques employed, systems based on UAD consistently exhibit reduced risks compared to those that do not incorporate this defense mechanism.
Figure 7: An outline of our study’s methodology. [57]
An AI-CAD model depicted in Figure 7 was initially trained to manipulate diagnosis-sensitive elements within images, such as adding or removing malignant tissue. The model’s effectiveness was then assessed using adversarial images produced by a GAN model. To determine how well human specialists could visually distinguish images produced by the GAN, a reader study was carried out.
In a study by Zhou et al. [58], the responses of an AI-CAD model to adversarial attacks on GAN-generated mammography images were investigated. This involved introducing malignant tissue into healthy images and removing cancerous tissue from images affected by cancer. The study also evaluated the proficiency of experienced radiologists in visually identifying such adversarial images, both before and after instructional intervention. The study cohort, provided by the University of Pittsburgh Medical Center, included 1284 women, and 4346 mammography images were collected. Among these, 366 patients had biopsy-proven malignant breast cancer, while 918 patients had breast cancer assessed as benign (including benign signs).
Following training, the researchers assessed the model’s classification accuracy concerning both the original genuine test data and its equivalent GAN-generated adversarial counterparts. Two GAN-trained U-Net23 models were employed to generate adversarial images, flipping labels in the test set. The high-resolution (1728 x 1408) photographs indicated the categorization impact of the AI-CAD model on the test data. The AI-CAD model achieved an AUC of 0.82 on the test set, consisting of 364 genuine negative samples and 74 real positive samples. The adversarial GAN-generated images (with flipped labels) in the test set yielded an AUC of 0.94. These AUC values suggested that the adversarial collection of images effectively deceived the AI-CAD model. Additionally, 59.5% (44 out of 74 cases) of genuine positive images with a classification accuracy threshold of 0.5 were correctly labeled as positive, whereas 95.5% (42 out of 44 cases) of adversarial samples created by GANs successfully tricked the system by being classified as negative.
Figure 8: The Mean Structural Similarity Index Measure (SSIM) was computed between the authentic images in the test sets and the adversarial instances generated using FGSM or PGD at different perturbation levels. The SSIM values displayed represent the average across two model architectures (Inception-v3 and Densenet-121) [59].
Figure 8 displays the mean SSIM values for FGSM and PGD assaults across all photos. The Supplementary Material includes the SSIM data for every model. As can be shown, the effects of the identical disturbance applied to various imaging modalities on human visual perceptibility with the observed SSIM varies. The radiology images most clearly showed adversarial disturbances, with = 0.02 producing an already apparent, but very modest perturbation. At the same perturbation level for the ophthalmology and pathology photos, perturbations were nearly undetectable and became apparent with larger epsilon values. The authors examine black-box adversarial attacks on deep learning in healthcare imaging. [59]. They investigate three medical imaging sectors where deep learning algorithms are vulnerable. In all three datasets, perturbations calculated by FGSM showed lower SSIM than those calculated using PGD. This is a predicted outcome given that PGD optimizes perturbations based on their size and influence on model predictions.We opted to document attacks in our subsequent analyses employing 0.02, as this represented the maximum perturbation level that remained effective across all applications and attack strategies. Additionally, it exhibited greater transferability than an epsilon value of 0.01 in the majority of the examined applications.
Figure 9: In a range of black-box settings, such as target and surrogate pre-trained on ImageNet, target and surrogate at random initialized, target and replacement randomly initialized plus surrogate developed using a different and reduced dataset (d2/2), FGSM (=0.02) was used to create the original images, adversarial images, and corresponding adversarial noise. The aforementioned shows the average receiver operating characteristic curve (AUC) for both clean and black-box scenarios for each configuration. Diabetic retinopathy (DR) is misclassified in a red frame and correctly classified as referable or non-referable in a green frame. The adversarial noise  represents the difference between the initial and adversarial image.citeadversarial.bortsova2021.
Figure 9 shows examples taken from the ophthalmology collection that demonstrate the transferability of assaults when the target and surrogate are randomly started and when both are pre-trained on ImageNet. Whether the target and surrogate models had the same architecture or a different one did not affect the findings.
5 Conclusion
Diagnostic imaging AI systems based on computer vision are increasingly being used as models for classifying and segmenting diseases. The scaled-up use of medical imaging AI systems has given rise to serious safety issues because to DNNs’ susceptibility to adversarial samples. Recently, a number of ways have been put out to increase the efficacy of medical image defense tactics. Although numerous protection strategies have been put forth, there are still reservations over the use of medical deep learning techniques. This outcome arises from certain constraints within medical imaging, including a scarcity of high-quality image datasets and labeled data in comparison to more abundant datasets available for natural images. The efficacy of adversarial defenses may not be universally applicable in this context. The findings underscore the importance of exercising caution in the design of Deep Neural Networks (DNNs) for medical imaging and their subsequent real-world applications.
Declarations
Funding
A portion of the funding for this research comes from DOEd Grant P116Z220008 (1) and NSF Grant No. 2306109. The author(s) expresses all opinions, findings, and conclusions; the sponsor(s) does not necessarily agree with them.
Conflict of interest
The authors do not declare any conflicts of interest.
Ethics approval
There is no original research involving human subjects, animal subjects, or sensitive data in this review publication.
Consent to participate
The studies and data included in this review were sourced from publicly available and previously published literature.
Consent for publication
All authors listed on this review paper have reviewed and approved the final manuscript for submission to Machine Learning. Each author has contributed significantly to the research, writing, and revision of the paper. All authors have read and agreed to the content presented in this review, and their consent for publication is hereby provided.
Availability of data and materials
The data and materials used can be found in the references of this work.
Code availability
The review is based on a comprehensive analysis of publicly available literature, and any references to specific methodologies or algorithms are attributed to the respective original sources. We have strived to provide clear citations and references to enable readers to locate the original works for further examination.
Authors’ contributions
Conceptualization: A.B., M.A.A.N.; Methodology: A.B., M.A.A.N.; Formal analysis and investigation: A.B., M.A.A.N.; Writing - original draft preparation: A.B., M.A.A.N., R.G., A.R.; Writing - review and editing: M.A.A.N., A.R.; Funding acquisition: K.D.G; Resources: R.G.; Supervision: K.D.G., R.G. All authors have read and agreed to the published version of the manuscript.

### v2

@@
 ML’s proponents laud its capacity to handle vast, complicated, and erratic medical data.
-It’s common knowledge that attackers might cause misclassification by deliberately creating inputs for machine learning classifiers.
-Research on adversarial examples has been extensively conducted in the field of computer vision applications.
+It is well known that adversaries can force misclassification by constructing inputs maliciously into machine learning classifiers.
+In the field of computer vision applications, such adversarial examples have been thoroughly researched.
 Healthcare systems are thought to be highly difficult because of the security and life-or-death considerations they include, and performance accuracy is very important.
 Recent arguments have suggested that adversarial attacks could be made against medical image analysis (MedIA) technologies because of the accompanying technology infrastructure and powerful financial incentives.
-Since the diagnosis will be the basis for important decisions, it is essential to assess how strong medical DNN tasks are against adversarial attacks.
+Since the diagnosis will be the basis for important decisions, assessing how strong medical DNN tasks are against adversarial attacks is essential.
 Simple adversarial attacks have been taken into account in several earlier studies.
@@
 keywords:
-Adversarial attack, Medical image, Deep Neural Network, Model safety, Robustness
+Adversarial attack; Medical image; Deep Neural Network; Model safety; Robustness
 1 Introduction
-1.1 Adversarial Attack
-Adversarial attacks encompass a set of techniques aimed at manipulating machine learning models by introducing well-crafted, often imperceptible alterations to input data.
-These modifications aim to deceive the model, leading to misclassifications or inaccurate predictions [1].
-Adversarial attacks have the potential to compromise machine learning systems’ dependability and security, which is concerning for crucial applications such as cybersecurity, autonomous vehicles, and medical diagnosis.
-Of all the adversarial attack types, ”White-box attacks” and ”Black-box attacks” are the most common.
-Ongoing research in machine learning security centers on the analysis of adversarial attacks and the development of defense strategies.
-Researchers continually explore novel attack tactics to identify potential weaknesses in AI systems while concurrently striving to construct robust models less susceptible to adversarial perturbations.
-Artificial intelligence (AI) modern technology is widely successful in fields like computer vision, natural language processing, and automated driving; however, its application in vital safety sectors is hindered by its susceptibility to adversarial attacks.
-Consequently, enhancing the resilience of AI systems against such attacks has become paramount for the progress of AI [2].
-The rapid advancement of AI technologies has found diverse applications in numerous domains, from machine translation, speech recognition, and object identification to more intricate tasks like drug composition analysis [3, 4, 5, 6, 7].
-Noteworthy applications include brain circuit construction [8], particle accelerator data analysis [9], [10], and DNA mutation impact analysis [11].
-Since Szegedy et al.’s seminal work [12] highlighting neural networks’ vulnerability to adversarial attacks, research on adversarial technologies for artificial intelligence has steadily expanded.
-New techniques for adversarial attacks and strategies for mitigating them are consistently emerging.
-The term ”adversarial attacks at the training stage” refers to actions taken by adversaries to manipulate the training dataset, input characteristics, or data labels during the training phase of the target model.
-This manipulation of the training dataset involves actions such as adding or deleting training data, as demonstrated by Barreno et al.’s approach [13].
-During the testing phase, adversarial assaults are categorized into white-box attacks and black-box attacks [2].
-In white-box situations, attackers have access to the parameters, techniques, and structure of the target model, enabling them to craft adversarial samples based on this information.
+In recent years, machine learning has been used as a transformative tool, particularly in medical image analysis.
+Particularly the processing ability of vast and complex medical data showed potential diagnostic accuracy and gave better outcomes.
+However, though there are advancements in machine learning security, there are several situations of adversarial attacks.
+These attacks, where inputs are maliciously altered to deceive the model, have been widely studied in computer vision but are less explored in the domain of medical imaging.
+This paper aims to explore recent adversarial attack strategies targeting DNNs in medical imaging, review detection techniques, and propose potential countermeasures to enhance the resilience and security of ML models in this critical field.
+We will go through the necessary sections to provide a critical review of these paper.
+The contributions of this research paper can be summarized below:
+  1. 1.
+    The research provides an in-depth overview of various adversarial attack strategies targeting deep neural networks (DNNs) in medical image analysis to identify vulnerabilities.
+  2. 2.
+    It examines existing DNN models used in medical image analysis and their resilience to adversarial threats.
+  3. 3.
+    The study evaluates the effectiveness of these DNN models against adversarial attacks.
+  4. 4.
+    Recommendations for enhancing security systems in medical image analysis are also provided.
+1.1 Adversarial Attacks ML Applications
+Adversarial attacks are a group of methods for altering machine learning models by adding expertly designed, frequently undetectable changes to input data.
+These alterations are intended to trick the model, causing it to misclassify data or forecast outcomes [1].
+Adversarial assaults can endanger the dependability and security of machine learning systems, especially in crucial applications like autonomous vehicles, medical diagnosis, and cybersecurity.
+A number of adversarial attack types exist, but the two most popular subtypes are ”White-box attacks” and ”Black-box attacks.” An continuing area of research in the realm of machine learning security is the analysis of adversarial assaults and defense techniques.
+Researchers are always exploring novel attack tactics to find possible weaknesses in AI systems while also attempting to create robust models that are less vulnerable to adversarial perturbations.
+Computer vision, natural language processing, automated driving, and other industries have all benefited greatly from the application of artificial intelligence technology in recent years [2].
+The uses of artificial intelligence (AI) technology in important security sectors are, however, constrained by the adversarial assaults that AI systems are susceptible to.
+Therefore, enhancing the resilience of AI systems against adversarial attacks has become more and more crucial to the advancement of AI [2].
+Artificial intelligence technology applications have recently advanced quickly in many different industries.
+Artificial intelligence technologies have been used in fields such as image classification, object detection, voice control, machine translation, and more advanced ones like drug composition analysis [3], brain circuit reconstruction [4], particle accelerator data analysis [5], [6], and DNA mutation impact analysis [7] due to their high performance, availability, and intelligence.
+Since Szegedy et al.’s [8] hypothesis that neural networks are susceptible to adversarial attacks, research on artificial intelligence adversarial technologies has steadily become a hotspot, with new adversarial attack techniques and mitigation strategies being developed on a regular basis.
+The term ”adversarial attacks at the training stage” refers to actions taken by the adversaries to alter the training dataset, input characteristics, or data labels during the training phase of the target model.
+By changing or removing training data, Barreno et al.’s [9] alteration of the training dataset falls under the category of training dataset modification.
+Adversarial attacks can be categorized into white-box attacks and black-box attacks during the testing stage [2].
+In white-box scenarios, the attackers have access to the target model’s parameters, methods, and structure.
+Adversaries can use this knowledge to create adversarial samples for attacks.
 1.2 The Adversary’s Objective: Evasion attack versus poisoning assault
-”Poisoning attacks” are assault methods that allow an attacker to add or change a large number of fake samples to the training set of a DNN algorithm.
+The term ”poisoning attacks” refers to the assaulting techniques that enable an attacker to insert/modify a number of fictitious samples into a DNN algorithm’s training database.
 The trained classifier may perform poorly as a result of these bogus data.
-They may have poor accuracy [14] conversely make imprecise predictions on a few samples for analysis [15].
+They may have poor accuracy [10] or make incorrect predictions on some test samples [11].
 The classifiers used in evasion attacks are fixed and often work well on safe testing samples.
-The adversaries are unable to reform the classifier’s settings or parameters, but they do generate some fictitious instances that the classifier is unable to distinguish.
+The adversaries are unable to alter the classifier’s settings or parameters, but they do create some fictitious samples that the classifier is unable to distinguish.
 1.3 Deep Neural Networks: Adversarial Attacks
-Regarding medical image processing use cases such as diagnosing cancer and lesion identification, deep neural networks, also known as DNNs, are rapidly gaining popularity [16].
-Nonetheless, a new study suggests that adversarial examples/attacks with tiny, barely noticeable changes can weaken medical deep learning systems.
-There are presently safety concerns associated with the usage of these devices in clinical settings [16].
-Deep neural networks (DNN) are increasingly well-liked and effective at a variety of machine learning requisitions.
-They have been employed with surprising effectiveness in a variety of recognition issues in the fields of pictures, graphs, text, and voice [1].
+Deep neural networks (DNNs) have gained popularity for medical image processing applications like cancer diagnosis and lesion detection [12].
+However, a recent study shows that adversarial examples/attacks with tiny, barely noticeable changes can weaken medical deep learning systems.
+The use of these devices in clinical settings is now accompanied by safety worries [12].
+Deep neural networks (DNN) are increasingly well-liked and effective at a variety of machine learning applications.
+They have been used with surprising effectiveness in a variety of recognition issues in the fields of pictures, graphs, text, and voice [1].
 They are able to identify things in images with accuracy that is almost human [13], [14].
@@
 Figure 1: The attack of Biggio’s SVM classifier for letter recognition. [1]
-Regarding the data set provided by MNIST, Biggio et al.[22] produce adversarial instances first, focusing on traditional machine learning classifiers like SVMs and 3-layer neural networks with full connectivity which is shown in Figure 1.
+On the MNIST data set, Biggio et al.[18] produce adversarial instances first, focusing on traditional machine learning classifiers like SVMs and 3-layer fully-connected neural networks which is shown in Figure 1.
 In order to trick the classifier, it optimizes the discriminant function.
@@
 The research [19] put decals to traffic signs that pose a serious threat to autonomous vehicles’ sign recognition technology.
-These hostile objects are especially detrimental to deep learning systems since they can interfere directly with numerous real-world DNN usages, such facial recognition and autonomous cars.
-By assessing the adversarial images (FGSM, BIM) produced to see if they are ”robust” against changes in natural circumstances (e.g., shifting viewpoint, illumination, etc.), the authors of the work [24] investigate the viability of creating tangible adversarial objects.
+These hostile objects can directly interfere with many real-world DNN applications, such as face recognition, driverless vehicles, etc., making them more harmful to deep learning models.
+By determining if the created adversarial images (FGSM, BIM) are ”robust” under natural change (such as changing viewpoint, illumination, etc.), the authors of the work [20] investigate the viability of creating tangible adversarial objects.
 Robust in this context means that even after transformation, the produced pictures are still antagonistic.
-The experiment’s findings show that many of these adversarial examples—particularly those produced by FGSM—remain antagonistic to the classifier even after transformation.
-The findings imply that antagonistic real-world objects could trick the sensor in many situations.
+The experimental findings show that a significant fraction of these adversarial examples, particularly those produced by FGSM, continue to be adversarial to the classifier following transformation.
+The findings point to the possibility of actual hostile objects that can trick the sensor in various settings.
 1.5 Medical Image Under Artificial Intelligence Attack
 Researchers have access to strong models of developing science and technology thanks to deep learning.
-Convolutional neural networks, also known as CNNs, are among the most significant categories of deep learning (DL) frameworks for the analysis and processing of images because of their exceptional ability to learn valuable information.
-The processing of the human organism utilizing different picture modalities for therapeutic, diagnostic, and health surveillance purposes is known as medical image analysis [25].
-The advancement of computer vision through the use of deep neural networks addresses issues that were not well-solved by traditional image processing methods.
+As they are highly good at learning useful features, convolutional neural networks (CNNs) are the most significant class of DL models for image processing and analysis.
+Medical image analysis is the processing of the human body using various picture modalities for diagnostic, therapeutic, and health monitoring purposes [21].
+The development of deep neural networks in the field of computer vision addresses issues that were not well-solved by traditional image processing methods.
 Beinfeld et al.
@@
 MRI, CT scans, ultrasound (US), and X-rays are the most frequently used image modalities.
-Figure 2 shows how adversarial assaults may be used to arbitrary alter diagnosis outcomes in three different medical picture datasets: Dermoscopy [27], X-ray of the chest [28], and Fundoscopy [29].
-Figure 2: Instances of adversarial approaches designed by the PGD (Projected Gradient Descent) to deceive DNNs developed using datasets of medical images include dermoscopy [27] (third row), chest x-ray [28], and fundoscopy [29] (first row, DR=diabetic retinopathy).
+Figure 1 illustrates how diagnosis outcomes can be arbitrarily changed by adversarial attacks across three medical picture datasets: Fundoscopy [23], Chest X-Ray [24], and Dermoscopy [25].
+Figure 2: Examples of adversarial approaches designed by the Projected Gradient Descent (PGD) to deceive DNNs trained on medical image datasets include dermoscopy [25] (third row), chest x-ray [24], and fundoscopy [23] (first row, DR=diabetic retinopathy).
 Normal images on the left, adversarial perturbations in the middle, and adversarial images on the right.
-The left-bottom tag indicates the projected class, and green or red indicates whether the predictions were accurate or not [16].
+The anticipated class is indicated by the left-bottom tag, and green or red denotes accurate or inaccurate predictions [12].
 Hu et al.
@@
 Depending on the knowledge of the enemy, attacks can be categorized into three kinds.
-The saliency (or attention) map of a picture that was input shows the regions that significantly change the model’s output based on the gradients of the categorization loss with regard to the input [16].
+Based on the gradients of the classification loss with respect to the input, the saliency (or attention) map of an input image indicates the regions that significantly alter the model’s output [12].
 We can see that some medical photos have highly concentrated regions that are noticeably larger.
-This could indicate that the rich biological textures included in medical photos occasionally entice the DNN model to focus more on aspects unrelated to the diagnosis.
-Tiny adjustments in these high-focus areas can significantly affect the model’s output.
+This could mean that the DNN model is occasionally drawn away by the rich biological textures present in medical images, causing it to focus more on details unrelated to the diagnosis.
+Small modifications in these areas of high focus can have a big impact on the model’s output.
 Due to several factors, including High-Dimensional Data, High-Dimensional Data, Transferability of Adversarial Examples, Complex Decision Boundaries, Black-Box Attacks, Safety-Critical Applications, and Safety-Critical Applications, Medical Image Deep Neural Network (DNN) models can be relatively easier to attack when compared to some other domains.
@@
 To further enhance the security and dependability of AI systems in healthcare applications, consistent evaluation frameworks for adversarial robustness in medical imaging must be developed.
-Byra et al.
-[30] proposed an attack strategy on ultrasonography (US) imaging for liver fatty tissue.
+An attack method on ultrasound (US) imaging for fatty liver was put forth by Byra et al.
+[26].
 Radio-frequency signals are used to rebuild US pictures, and the reconstruction technique was subjected to a zeroth-order optimization attack [27].
 The InceptionResNetV2 model was used in the studies, and the assault resulted in a 48% loss in model accuracy.
-Ozbulak et al.
-[32] proposed an attack specifically designed for medical picture segmentation called the adaptive segmentation masking attack (ASMA).
+Adaptive segmentation mask attack (ASMA), a specific attack for medical picture segmentation, was suggested by Ozbulak et al.
+[28].
 The suggested attack offers significant intersection-over-union (IoU) degradation and produces nearly undetectable samples for most portions.
-Because the U-Net framework is among the most well-known models for clinical picture segmentation, they employed it in the trials.
-The datasets used were for segmentation of glaucoma optic disk [33] and ISIC skin lesion segmentation [34].
-Chen et al.
-published a method for fabricating hostile cases to thwart medical picture segmentation [35].
+Because the U-Net model is one of the most well-known models for medical picture segmentation, they employed it in the trials.
+Datasets for ISIC skin lesion segmentation [29] and glaucoma optic disk segmentation [30] were employed.
+A technique for creating hostile instances to undermine medical picture segmentation was put out by Chen et al.
+[31].
 In order to simulate anatomical and intensity fluctuations, geometrical deformations are used to create the adversarial examples.
 By attempting to partition organs from abdominal CT scans using a U-Net model, they tested the effectiveness of these examples.
-With reference to the Dice score measure, they successfully reduced it significantly across all organs.
-The kidneys and pancreas, however, require a higher level of disturbance and are more arduous to assault than the liver as well as spleen.
+In terms of the Dice score metric, they successfully reduced it significantly across all organs.
+The kidneys and pancreas, however, require a higher level of disturbance and are more difficult to assault than the liver and spleen.
 Figure 3: Clean image, bias field noise, and diagnosis following application of bias field noise are shown in (a), (b), and (c), respectively [32].
-Tian et al.
-[36] examined the phenomena of bias field, which can arise from incorrectly capturing a medical image and jeopardize a DNN’s efficacy, as shown in Figure 3.
-Motivated by adversarial attacks, the authors created an adversarial-smooth bias discipline approach to trick a model.
-The ResNet50, MobileNet, among others and DenseNet121 models were employed to fine-tune the chest X-ray dataset that was used in their study.
+As seen in Figure 3, Tian et al.
+[32] looked at the phenomenon of bias field, which can result from improperly acquiring a medical image and compromise a DNN’s effectiveness.
+The authors developed an adversarial-smooth bias field attack to deceive a model after being inspired by adversarial assaults.
+The chest X-ray dataset used for their research was tuned using the ResNet50, MobileNet, and DenseNet121 models.
 They looked at this attack’s transferability and white-box attacks.
@@
 2 Background Literature Review
-Adversarial instances are intentionally created data inputs that are intended to degrade a machine learning model’s performance.
-When researchers examined strategies used by spammers to evade spam filters in 2004, they unofficially coined the term ”adversarial inputs” [37].
-These misleading examples are usually produced by deliberately faking real data, such spam advertising messages, in order to trick the computer system that analyzes it.
-Alterations can be applied to text data, like spam, by introducing innocuous text or modifying frequently used phrases in malicious communications with synonyms.
-To fortify algorithms against adversarial attacks, researchers have explored various strategies, including training algorithms on adversarial samples and employing sophisticated data processing techniques to minimize the susceptibility to manipulation.
-The quest for fully robust models in machine learning aims to accelerate the development of algorithms capable of making decisions based on consistent explanations, with promising early efforts in this direction [38].
-In the realm of healthcare, medical claims codes play a crucial role in determining the amount spent on a patient visit post-payer approval.
-Payers often assess these claims using automated fraud detectors increasingly driven by machine learning.
-Historically, healthcare providers have shaped payer records of patient visits, including the associated codes, to influence the algorithmic outputs of payers [38].
-Medical fraud, a market valued at \$250 billion, exemplifies the extreme end of this strategic tailoring of patient presentations.
-While some practitioners may overtly fabricate medical claims, patient data falsification often takes more covert forms.
-For instance, consistently submitting codes for billing services identical to, but more expensive than, those actually provided is known as intentional upcoding.
-Here, Table 1 shows past exclusive past research works where authors did some fantastic work for medical images to defend adversarial attacks.
-Different researcher proposed different frameworks to overcome this problem and to suggest a new thought for revolution.
-2.1 Elements of a Cyber Adversary’s Assault
-Figure 4: A showcase of executing adversarial attacks on various medical AI systems without resorting to overtly deceptive data manipulation [39].
-In some instances, subtle adjustments to billing codes blur the boundary between fraudulent activities and conscientious best practices.
-One notable illustration may be found in the guidelines on the Endocrine Society website, which advise medical professionals not to bill for metabolic syndrome, or International Classification of Diseases (ICD) code 277.77, in situations of obesity.
-Coverage denial is likely to occur with this particular code/condition combo in Figure 4.
-The Society recommends an alternative approach, suggesting billing for the individual codes related to specific disorders comprising the metabolic syndrome, such as hypertension.
-Table 1: Existing adversarial attack research on medical Image
-Reference Year Framework Contribution
-[40] 2021 Propose a framework to defend against adversarial training, speckle-noise attacks, and maintain accurate classification labels in recognizing diabetic retinopathy through analyzing retinal fundus images.
-a) accuracy of 99%; b) defensive model’s robustness ; c) a system that includes a fresh SN attack.
-[41] 2022 Explain the use of gradient-free trained sign activation systems in medical imaging AI systems to identify and thwart hostile attacks.
-The model performs better and even twice as well as the finest.
-a) Owing to significant distortion and excellent transferability accuracy, the model’s average success rate in adversarial case classification is 88.89%; b) This model defends against adversarial attacks.
-[42] 2020 This method detects diverse adversarial attacks while preserving user anonymity and classification performance, making it a versatile addition for deep learning-based medical imaging systems to enhance resilience.
-a) The chest X-ray dataset exhibits robust performance across diverse configurations, contributing to the heightened safety of deep learning-based medical image classification systems; b) To improve medical image categorization systems based on deep learning security.
-[43] 2022 In weakly-supervised clinical tasks, CNNs and ViTs exhibit vulnerability to both white- and black-box adversarial attacks with comparable baseline performance.
-a)Compared to CNNs, aViTs have a larger latent portrayal of clinically meaningful categories; b) Findings align with prior theoretical investigations and provide tangible evidence of ViTs’ capacity to grasp computational pathology; c) This suggests that computational pathology AI models will be deployed widely.
-[…]
+Adversarial examples are data inputs that are specifically designed to make a machine learning model fall short.
+The term ”adversarial inputs” was first used informally in 2004, when academics looked into the strategies spammers used to get around spam filters [33].
+Typically, threatening examples are created by intentionally altering genuine data, such as spam advertising messages, in order to deceive the algorithm that will process it.
+Such modifications can be made to text data, such as spam, by adding innocent text or by changing phrases that are frequently used in malicious communications to synonyms.
+Researchers have worked to create algorithms that are resistant to adversarial attacks, for example by exposing algorithms to hostile examples during training or utilizing cunning data processing to reduce the possibility of manipulation.
+Early efforts in this direction are encouraging, and we are hopeful that the pursuit of completely robust machine-learning models will spur the creation of algorithms that learn to make judgments for consistent justifications [34].
+Following payer approval, medical claims codes determine the amount paid for a patient visit.
+Payers often use automated fraud detectors, which are increasingly driven by machine learning, to assess these claims.
+Health care providers have historically shaped payers’ records of patient visits (and the related codes) to affect their decisions (the algorithmic outputs) [34].
+Medical fraud, a \$250 billion market, is at the extremity of this tactical tailoring of a patient presentation.
+Although some practitioners might blatantly fabricate medical claims, patient data falsification frequently takes much more covert forms.
+Intentional upcoding, for instance, is the practice of routinely submitting billing codes for services that are similar to but more expensive than those that were actually rendered.
 2.1 The components of an adversarial attack
-Figure 5: Distribution of chosen articles on a yearly basis Pie chart; line graph [44].
-Presenting a thorough overview of the various adversarial attack strategies and defense techniques is the aim of the study [44].
-We first discuss the theoretical underpinnings, practices, and practical uses of adversarial attack strategies.
-Then, a brief discussion of the research on defense strategies obscuring the boundary of the large field follows.
-A few selected articles are published year-by-year from 2012 to 2021, as seen in Figure 5.
-This work aims to provide thorough taxonomies, evaluations of harmful assaults, and protections against the full DL pipeline.
-In this context, the numerous attack and defense tactics that have been developed over the previous two years have been categorized, with a focus on the clinical deep learning systems that are susceptible to adversarial attacks.
-Several defensive strategies have been put forth in an effort to neutralize possible threats.
-A popular method in natural imaging called adversarial training adds adversarial images to the training dataset in order to make Convolutional neural network, or CNN, models more resilient.
-Nevertheless, this method is not the best for datasets pertaining to medical imaging, since the addition of different adversarial images could considerably reduce classification accuracy.
-A robust detection technique for malicious images is presented in the work of Li et al.
-[42], successfully defending against attacks on deep learning-driven medical image labeling systems.
-In a different study [43], scientists compared CNN performance with Vision Transformers (ViTs) to determine how resilient CNNs are to different types of attacks in computational pathology.
-The authors developed robust neural network models, evaluating their efficacy against both white- and black-box attacks.
-The structures of attacks for both models were scrutinized, with an exploration of the underlying factors influencing their performance.
-The study’s findings were validated through two clinically relevant classification tasks involving distinct patient groups.
-Preceding Ma et al.’s research [16], ambiguity surrounded medical image adversarial attacks (AAs), limiting adversarial machine learning analyses to natural images.
-Unlike natural images, medical images may contain domain-specific elements, impacting medical deep-learning systems susceptible to AAs.
-AAs possess the potential to manipulate diagnoses and outcomes, emphasizing the critical need for a robust healthcare infrastructure to mitigate potential risks from malicious attacks.
+The goal of the study [35] is to present a comprehensive overview of the various adversarial assault strategies and defense techniques.
+First, we talk about the theoretical foundations, procedures, and applications of adversarial attack tactics.
+The research on defense strategies covering the large field’s boundary is then briefly discussed.
+Figure LABEL:fig5 shows the year-by-year publishing of a few chosen articles from 2012 to 2021.
+In-depth taxonomies and analyses of harmful attacks and defenses against the full DL pipeline are the goals of this research.
+The numerous existing attack and defense strategies were categorized in this context, with a focus on those that have only been developed in the last two years, and the medical deep learning systems sensitive to adversarial attacks were examined.
+Numerous defensive strategies have been suggested as a defense against the threat.
+Adversarial training, which enlarges the training dataset with adversarial images to improve the resilience of the trained Convolutional Neural Network (CNN) model, is a common technique in the natural imaging domain.
+This tactic is not ideal for medical imaging datasets, though, as the introduction of a large number of different hostile images into the training set can seriously reduce the classification accuracy.
+In article [36], authors suggest a reliable detection method for malicious images that can successfully fend off attacks on deep learning-based medical image categorization systems.
+In the study [37], scientists looked at how resilient CNNs in computational pathology were to various attacks and compared the findings to how resilient ViTs were.
+Authors also developed strong neural network models and assessed how well they defended against white- and black-box attacks.
+The attack structures for both models were examined, and the authors looked into why they performed as they did.
+The findings were verified by the authors using two clinically pertinent classification tasks on separate patient groups.
+Prior to the publication of Ma et al.’s study [12], medical image AAs were ambiguous and adversarial machine learning analysis was limited to natural image analysis.
+Medical images, in contrast to natural images, may include domain-specific elements.
+Medical deep-learning systems may be impacted by AAs.
+AAs have the discretion to alter diagnoses and results.
+Ma et al.
+emphasized the significance of the healthcare system’s extensive resource base.
+It invariably raises dangers where potential attackers might try to take advantage of the healthcare system.
 3 Materials and Methods
+This study examines various strategies to strengthen the usability of medical image analysis models against adversarial attacks.
+These strategies include techniques such as input denoising, adversarial training, adversary detection, image-level pre-processing, feature enhancement, and knowledge distillation.
+Each method tries to improve the aspects of models, which is necessary to the potential risks these attacks pose to clinical applications.
+Additionally, the research presents a new classification system for adversarial defense in medical imaging, categorized by application type, covering white-box, semi-white-box, and black-box attacks.
+By addressing the vulnerabilities in computerized diagnostic systems, this work contributes to the development of reliable and accurate healthcare technologies, ensuring consistent performance even under adversarial conditions.
 3.1 Medical Adversarial Defensive Strategies
-Various defense models, including techniques such as input denoising, input gradients regularization, and adversarial training, have been devised.
-However, recent attacks often manage to evade or partially circumvent these protective measures.
-This paper undertakes a comprehensive exploration of adversarial attack and defense methods within the domain of medical image analysis.
-It introduces a novel taxonomy based on the application context for a family of techniques, as outlined in Dong et al.’s work [45].
-The author establishes a unified theoretical framework encompassing diverse adversarial attack and defense strategies tailored for medical imaging applications.
-The primary emphasis in most research efforts on adversarial attacks in medical image processing is on the white-box scenario.
-These studies, characterized by a comprehensive understanding of medical Deep Neural Networks (DNNs), concentrate on the susceptibility of computer-aided diagnosis models across various medical imaging applications.
-In the execution of adversarial attacks, attackers may utilize the target diagnosis DNN as a locally deployed model.
-In addition to white-box adversarial attacks for medical classification tasks, scholars have delved into vulnerabilities associated with various medical imaging tasks.
-Notably, these publications focus predominantly on medical segmentation, as evident in works such as Chen et al.’s [35] and Ozbulak et al.’s [32].
+Defense models such as input denoising, input gradients regularization, and adversarial training have been created.
+The most recent attacks can typically completely or partially escape these countermeasures.
+In this case, a thorough investigation of adversarial attack and defense techniques in the field of medical image analysis is presented, including a family of techniques with a novel taxonomy based on the application situation [38].
+A unified theoretical framework was built by the author for various adversarial attack and defense strategies in medical imaging applications.
+The white-box scenario is the main focus of most efforts on adversarial attack for medical image processing.
+These studies, in particular, with complete understanding of medical DNNs, focus on the adversarial vulnerability of computer-aided diagnosis models in diverse medical imaging applications.
+When executing adversary creation, the attacker might use the target diagnosis DNN as a locally deployed model.Additional academics have looked into the weaknesses of additional medical imaging tasks in addition to white-box adversarial attacks for medical classification tasks.
+These publications mostly focus on medical segmentation [31] [28].
 For natural photos, semi-white-box (Gray-box) attacks have been extensively researched [39].
@@
 2) Instead of needing to know anything about the target model, as would be the case in a completely black-box scenario, the adversary generator can directly obtain adversarial examples against the target model with the input of authentic photos during the application stage.
-Presently utilized white-box adversarial attacks often require multiple backward gradients of the target model.
-In simpler terms, attackers generate comparable adversarial samples by treating the target Deep Neural Network (DNN) as if it were a locally deployed model.
-However, due to its reliance on an in-depth understanding of the DNN model to execute an attack, the white-box approach may be unreliable in real-world scenarios.
-Conversely, the general black-box scenario may offer a more suitable environment for simulating real-world adversarial attacks, with numerous proposals focusing on investigating black-box assaults for natural images.
-The outlined solutions usually involve multiple queries to the black-box model or depend on having a comprehensive understanding of the desired diagnosis model.
-However, in many real-world situations, attackers may lack direct access to the target diagnosis model.
-The restricted black-box (no-box) configuration, specifically representing the most challenging scenario for real-world adversarial assaults, may be more covertly dangerous even without querying the target black-box DNN.
-The transferability [49] of adversarial instances across different DNN models fundamentally determines the success of no-box attacks.
-For example, a no-box attacker can generate adversarial images based on a locally deployed surrogate model, transferring them directly to target medical diagnosis systems.
-Restricted black-box adversarial attacks are deemed more subtly perilous for natural vision tasks according to a renowned study.
-Considering the substantial risks to the healthcare sector, numerous defense strategies have been proposed to counter medical adversarial attacks.
-In order to overcome the weaknesses shown by adversarial examples, a number of methods have been developed to provide trustworthy deep learning-based systems for natural pictures [49].
-The development of accurate computer-aided diagnosis models for clinical applications not only contributes to trustworthy healthcare services for millions but also underscores the importance of investigating adversarially robust models within the realm of medical image analysis.
-In their work [45], the authors succinctly summarize studies on adversarial defense in the context of medical image analysis.
-Various attack methods in the realm of medical images serve as robustness evaluation criteria for adversarial defense, complementing the extensive range of adversarial attack methods mentioned.
-The minimal separation between adversarial examples and the model’s decision boundary signifies the model’s resistance to disturbances.
-The goal of adversarial training is to mitigate subsequent losses.
-Minimize the
-$L(f(x_{\text{adv}},y)+\lambda\cdot L(x,x_{\text{adv}}))$ (1)
-where $L(f(x_{\text{adv}},y)$ is the classification loss, $L(x,x_{\text{adv}})$ is the perturbation size, and $\lambda$ is a trade-off hyperparameter.
-Transforming the input data into an alternative feature space that diminishes the impact of adversarial perturbations on the model.
-A technique aimed at enhancing resilience against adversarial attacks involves introducing noise to the features or intermediate representations that characterize the input data.
+Currently used white-box adversarial attacks mostly call for numerous target model backward gradients.
+In other words, the attacker generates equivalent adversarial samples by treating the target DNN as the locally deployed model.
+However, because it requires a comprehensive understanding of the DNN model to attack, the white-box option can be unreliable in a real-world scenario.
+Comparatively, the general black-box scenario may be a better environment for simulating real-world adversarial attacks.
+There have been numerous proposals to investigate black-box assaults for natural images.
+The solutions outlined above either require numerous queries to the black-box model or depend on having complete understanding of the desired diagnosis model.
+However, in the majority of real-world scenarios, the attacker might not be able to directly access the target diagnosis model.
+The restricted black-box (no-box) configuration in particular might effectively represent the toughest (worst) situation for the real-world adversarial assault, even if it doesn’t call for querying the target blackbox DNN.
+The transferability [42] of adversarial cases across different DNN models is the fundamental determinant of the no-box attack.
+For instance, a no-box attacker can create adversarial images based on a surrogate model that has been deployed locally and that can be transferred straight to the target medical diagnosis systems.
+Restricted black-box adversarial attacks are more covertly dangerous for natural vision tasks, according to a well-known study.
+Given the significant dangers to the healthcare sector, a number of defense strategies have been put out to fend off medical hostile attacks.
+Numerous approaches have been put out to provide reliable deep learning-based systems for natural photos in response to the catastrophic failures brought on by adversarial instances [42].
+Developing accurate computer-aided diagnosis models for clinical applications also helps millions of people receive trustworthy healthcare services.
+Investigating adversarially robust models in the context of medical image analysis is therefore important.
+In [38], authors summarize the works on adversarial defense in the context of medical imagine analysis.
+In the context of medical pictures, various attack methods also serve as robustness evaluation criteria for adversarial defense, in addition to a considerable number of adversarial attack methods stated here.
 3.2 Adversarial Training
@@
 3.3 Adversarial Detection
-Adversary detection seeks to identify adversary cases from input examples during the application stage, as opposed to developing robustness during the training stage of computer-aided diagnosis models.Many adversarial detection techniques have been put forth in the field of medical image analysis to stop additional misdiagnosis caused by adversarial samples [42].
+Adversary detection seeks to identify adversary cases from input examples during the application stage, as opposed to developing robustness during the training stage of computer-aided diagnosis models.
+In the field of biological image analysis, a variety of adversarial detection techniques have been put forth to prevent further misdiagnosis brought on by adversarial examples [36].
 In particular, it is possible to think about medical adversarial detection as an anomaly detection issue that can be resolved by combining explainability approaches [46].
@@
 Image-level pre-processing can be useful and secure in the context of biomedical image analysis because it does not require re-training or modifying medical models.
-3.5 Improvement of Features
-The difference in robustness between human and machine vision is attributed to adversarial examples linked to non-robust features extracted from specific patterns in the data distribution [54].
-Consequently, enhancing feature representation is crucial for the development of robust inference systems.
-In this investigation, we characterize feature augmentation as the modification of architectures or mapping functions.
-Various techniques aimed at improving features have significantly enhanced the robustness of medical classification models [55].
+3.5 Enhancement of Features
+The discrepancy between the robustness of human and machine vision has been shown to be caused by adversarial examples being associated with non-robust features (extracted from specific patterns in the data distribution) [47].
+Therefore, improving the feature representation is absolutely required for robust inference systems.
+In this study, we define the feature augmentation as the alteration of architectures or mapping functions.
+The robustness of medical classification models has been increased by the development of numerous feature improvement techniques [48].
 3.6 Distillation of Knowledge
-In the field of machine learning, knowledge distillation is a helpful technique for transferring learned information from a complex (teacher) model to a simple (student) model.
-Therefore, the situation where the network structures for the teacher and learners are identical is specifically referred to as self-distillation.
-Furthermore, a great deal of research has been done on the natural imaging domain using adversarial knowledge distillation, which moves the adversarial resilience from a heavy teacher model to a view student model [56].
+Knowledge distillation is a useful method for moving learnt information from a complicated (teacher) model to a simple (student) model in the machine learning field.
+As a result, self-distillation refers particularly to the scenario in which the network architectures for the teacher and student models are the same.
+Additionally, adversarial knowledge distillation, which transfers the adversarial resilience from the heavy teacher model to a light student model, has been extensively investigated for the natural imaging domain [49].
 4 Result and Discussion
-This study utilizes four publicly accessible benchmark datasets, as detailed in [45], to investigate adversarial attack and defense in the context of medical image processing (illustrated in Figure 6).
-Firstly, the Messidor1 dataset comprises 1,200 eye fundus color images for detecting diabetic retinopathy across four classes based on retinopathy grade.
-Secondly, the International Skin Imaging Collaboration (ISIC) dataset consists of 2,750 dermoscopic images categorized into three classes for skin lesion classification and segmentation.
-Thirdly, the ChestX-ray 14 dataset comprises 112,120 frontal-view X-ray images representing 14 thorax diseases.
-Lastly, the COVID-19 database incorporates 21,165 chest X-ray images accompanied by segmentation-capable lung masks.
+Four publicly available benchmark datasets are used in this study [38] to explore adversarial attack and defense for medical image processing (shown in Figure 4).
+2) Messidor1 dataset of 1,200 eye fundus color numerical pictures for detecting diabetic retinopathy of four classes according to retinopathy grade; International Skin Imaging Collaboration (ISIC) dataset of 2,750 dermoscopic images of three different categories for skin lesion classification and segmentation.
+3) The ChestX-ray 14 dataset consists of 112,120 frontalview X-ray images from 14 thorax diseases.
+4) The COVID-19 database, which includes 21,165 chest X-ray images with segmentation-capable lung masks.
 Figure 4: Medical adversarial examples with predictions for a range of perturbation sizes. For visualization, the created segmentation masks are placed on top of the source photos [38].
-The study conducted by the authors [57] employs experiments to demonstrate the following: 1) The SSAT module significantly enhances the adversarial robustness of the model without compromising the classification accuracy of clean images.
-2) The UAD module effectively identifies and excludes a majority of successful adversarial examples.
-3) In comparison to other existing AI systems, their medical imaging AI solution (UAD + SSAT) minimizes the risk of adversarial attacks.
-The research utilizes a publicly available dataset of retinal OCT images, employing the ”white-box” setting as the most challenging threat for assessing class prediction performance.
-Across all white-box attack conditions, the authors establish that SSAT consistently outperforms alternative baselines while maintaining comparable or superior performance for clean image classification.
-Furthermore, the authors demonstrate that the combined approach of UAD complementing SSAT results in the lowest adversarial risk according to the new measure presented.
-Regardless of the training techniques employed, systems based on UAD consistently exhibit reduced risks compared to those that do not incorporate this defense mechanism.
+Experiments are used by authors [50] to show: 1): Without affecting the classification accuracy of clean images, the SSAT module can considerably improve the adversarial robustness of the model.
+2) The bulk of successful adversarial examples may be found and excluded by the UAD module.
+3) When compared to other existing AI systems, their medical imaging AI solution (UAD + SSAT) minimizes adversarial risk.
+A publicly available dataset of retinal OCT images is used for the research.
+A publicly available dataset of retinal OCT images is used for the research.
+The most difficult threat used to assess class prediction performance is the ”white-box” setting.In all white-box attack conditions, the authors show that SSAT significantly outperforms alternative baselines while maintaining an equivalent or superior performance for clean picture classification.
+The authors show that the lowest adversarial risk for the new measure presented results from UAD complementing with SSAT.
+Regardless of the training techniques employed, it is evident that UAD-based systems consistently have reduced risks compared to those who do not.
 Figure 5: An outline of our study’s methodology. [50]
-An AI-CAD model depicted in Figure 7 was initially trained to manipulate diagnosis-sensitive elements within images, such as adding or removing malignant tissue.
-The model’s effectiveness was then assessed using adversarial images produced by a GAN model.
-To determine how well human specialists could visually distinguish images produced by the GAN, a reader study was carried out.
-In a study by Zhou et al.
-[58], the responses of an AI-CAD model to adversarial attacks on GAN-generated mammography images were investigated.
-This involved introducing malignant tissue into healthy images and removing cancerous tissue from images affected by cancer.
-The study also evaluated the proficiency of experienced radiologists in visually identifying such adversarial images, both before and after instructional intervention.
-The study cohort, provided by the University of Pittsburgh Medical Center, included 1284 women, and 4346 mammography images were collected.
-Among these, 366 patients had biopsy-proven malignant breast cancer, while 918 patients had breast cancer assessed as benign (including benign signs).
-Following training, the researchers assessed the model’s classification accuracy concerning both the original genuine test data and its equivalent GAN-generated adversarial counterparts.
-Two GAN-trained U-Net23 models were employed to generate adversarial images, flipping labels in the test set.
-The high-resolution (1728 x 1408) photographs indicated the categorization impact of the AI-CAD model on the test data.
-The AI-CAD model achieved an AUC of 0.82 on the test set, consisting of 364 genuine negative samples and 74 real positive samples.
-The adversarial GAN-generated images (with flipped labels) in the test set yielded an AUC of 0.94.
-These AUC values suggested that the adversarial collection of images effectively deceived the AI-CAD model.
-Additionally, 59.5% (44 out of 74 cases) of genuine positive images with a classification accuracy threshold of 0.5 were correctly labeled as positive, whereas 95.5% (42 out of 44 cases) of adversarial samples created by GANs successfully tricked the system by being classified as negative.
-Figure 8: The Mean Structural Similarity Index Measure (SSIM) was computed between the authentic images in the test sets and the adversarial instances generated using FGSM or PGD at different perturbation levels.
-The SSIM values displayed represent the average across two model architectures (Inception-v3 and Densenet-121) [59].
+An AI-CAD model, in Figure 5, that intended to alter the diagnosis-sensitive contents of images (by adding or deleting malignant tissue) was first trained, and then it was evaluated on the adversarial images produced by the GAN model.
+The reader study looked at how well human specialists could visually identify the images created by the GAN.
+Authors [51] of a research paper conducted a study to assess an AI-CAD model’s responses to adversarial attacks on GAN-generated mammography pictures by introducing malignant tissue into healthy photos and removing cancerous tissue from cancer-affected images.
+Additionally, they assessed the effectiveness of experienced radiologists in recognizing these types of antagonistic pictures visually both without and after an instructional intervention.
+The University of Pittsburgh Medical Center provided the authors with 1284 women for their study cohort, and they also collected 4346 mammography pictures for this cohort.
+In this cohort, 366 patients had breast cancer that was biopsy-proven to be malignant while 918 patients had breast cancer that was evaluated as benign (including benign signs).
+Following training, researchers evaluated the model’s classification accuracy in comparison to both the original genuine test data and its equivalent GAN-generated adversarial counterparts.
+To generate the adversarial images, we created two GAN-trained U-Net23 models.
+The GAN generators, which were trained to create fake positive mammography images from the falsely negative mammogram photos and fake negative mammogram images from the falsely positive mammogram images, respectively, updated the mammogram images in the test set with flipped labels.
+The two distinct resolutions of this image indicate the categorization impacts of the AI-CAD model on the test data.
+The authors begin by outlining the findings from the high-resolution (1728 1408) photographs.
+As can be observed, the AI-CAD model obtained an AUC of 0.82 on the test set, which consisted of 364 genuine negative samples and 74 real positive samples.
+The test set’s equivalent adversarial GAN-generated images (with flipped labels) yielded an AUC of 0.94.
+These two AUC values show that the adversarial collection of photos generally succeeded in deceiving the AI-CAD model.
+Additionally, when calculating classification accuracy using a threshold of 0.5, 44 out of the 74 real positive images (or 59.5%) were correctly labeled as positive cases; however, only 42 out of the 44 fake negative images (note that their labels were negative due to the flipping) were categorized as negative cases, indicating that 95.5% (42 out of 44 cases) of the GAN-generated adversarial samples successfully deceived the system.
+Figure 6: Between the original images in the test sets and the adversarial examples produced using FGSM or PGD with varied perturbation degrees, the Mean Structural Similarity Index Measure (SSIM) was calculated.
+The SSIM values shown by the points are averaged over two model architectures (Inception-v3 and Densenet-121).[52]
 Figure 6 displays the mean SSIM values for FGSM and PGD assaults across all photos.
@@
 At the same perturbation level for the ophthalmology and pathology photos, perturbations were nearly undetectable and became apparent with larger epsilon values.
-The authors examine black-box adversarial attacks on deep learning in healthcare imaging.
-[59].
+Black-box adversarial assaults on deep learning in medical imaging are studied by the authors [52].
 They investigate three medical imaging sectors where deep learning algorithms are vulnerable.
 In all three datasets, perturbations calculated by FGSM showed lower SSIM than those calculated using PGD.
-This is a predicted outcome given that PGD optimizes perturbations based on their size and influence on model predictions.We opted to document attacks in our subsequent analyses employing 0.02, as this represented the maximum perturbation level that remained effective across all applications and attack strategies.
-Additionally, it exhibited greater transferability than an epsilon value of 0.01 in the majority of the examined applications.
-Figure 9: In a range of black-box settings, such as target and surrogate pre-trained on ImageNet, target and surrogate at random initialized, target and replacement randomly initialized plus surrogate developed using a different and reduced dataset (d2/2), FGSM (=0.02) was used to create the original images, adversarial images, and corresponding adversarial noise.
-The aforementioned shows the average receiver operating characteristic curve (AUC) for both clean and black-box scenarios for each configuration.
-Diabetic retinopathy (DR) is misclassified in a red frame and correctly classified as referable or non-referable in a green frame.
-The adversarial noise  represents the difference between the initial and adversarial image.citeadversarial.bortsova2021.
-Figure 9 shows examples taken from the ophthalmology collection that demonstrate the transferability of assaults when the target and surrogate are randomly started and when both are pre-trained on ImageNet.
-Whether the target and surrogate models had the same architecture or a different one did not affect the findings.
-5 Conclusion
+This is a predicted outcome given that PGD optimizes perturbations based on their size and influence on model predictions.We decided to report attacks using for our further investigations utilizing = 0.02, since this was the maximum degree of disturbance that was still for all applications and assault strategies, and it had more transferability than an epsilon of 0.01 in most cases of the applications examined.
+Figure 7: FGSM (=0.02) was used to create the original images, adversarial images, and corresponding adversarial noise in a variety of black-box settings, including target and surrogate pre-trained on ImageNet, target and surrogate randomly initialized, target and surrogate randomly initialized plus surrogate developed using a different and reduced dataset (d2/2).
+The average receiver operating characteristic curve (AUC) for each configuration for both clean and black-box scenarios is shown above.
+Diabetic retinopathy (DR) is correctly classified as referable or non-referable in a green frame and incorrectly classified in a red frame.
+The difference between the original and the adversarial image is represented by the adversarial noise [52].
+Examples from the ophthalmology dataset are shown in Figure 7 to demonstrate attack transferability when the target and surrogate are both pre-trained on ImageNet and when both are initialized randomly.
+The instances where the target and surrogate models had the same or different architectures had no effect on any of the aforementioned results.
+5 Conclusion and Discussion
+The study identifies the impact of adversarial defense strategies in medical image analysis using datasets like Messidor1, ISIC, ChestX-ray14, and COVID-19.
+The SSAT module significantly improves model robustness against adversarial attacks while maintaining accuracy on clean images.
+Combined with the UAD module, it efficiently detects and removes adversarial examples, minimizing risks in upcoming applications.
 Diagnostic imaging AI systems based on computer vision are increasingly being used as models for classifying and segmenting diseases.
@@
 Although numerous protection strategies have been put forth, there are still reservations over the use of medical deep learning techniques.
-This outcome arises from certain constraints within medical imaging, including a scarcity of high-quality image datasets and labeled data in comparison to more abundant datasets available for natural images.
-The efficacy of adversarial defenses may not be universally applicable in this context.
-The findings underscore the importance of exercising caution in the design of Deep Neural Networks (DNNs) for medical imaging and their subsequent real-world applications.
-Declarations
-Funding
-A portion of the funding for this research comes from DOEd Grant P116Z220008 (1) and NSF Grant No.
-2306109.
-The author(s) expresses all opinions, findings, and conclusions; the sponsor(s) does not necessarily agree with them.
-Conflict of interest
-The authors do not declare any conflicts of interest.
-Ethics approval
-There is no original research involving human subjects, animal subjects, or sensitive data in this review publication.
-Consent to participate
-The studies and data included in this review were sourced from publicly available and previously published literature.
-Consent for publication
-All authors listed on this review paper have reviewed and approved the final manuscript for submission to Machine Learning.
-Each author has contributed significantly to the research, writing, and revision of the paper.
-All authors have read and agreed to the content presented in this review, and their consent for publication is hereby provided.
-Availability of data and materials
-The data and materials used can be found in the references of this work.
-Code availability
-The review is based on a comprehensive analysis of publicly available literature, and any references to specific methodologies or algorithms are attributed to the respective original sources.
-We have strived to provide clear citations and references to enable readers to locate the original works for further examination.
-Authors’ contributions
-Conceptualization: A.B., M.A.A.N.; Methodology: A.B., M.A.A.N.; Formal analysis and investigation: A.B., M.A.A.N.; Writing - original draft preparation: A.B., M.A.A.N., R.G., A.R.; Writing - review and editing: M.A.A.N., A.R.; Funding acquisition: K.D.G; Resources: R.G.; Supervision: K.D.G., R.G.
-All authors have read and agreed to the published version of the manuscript.
+This is a result of some limitations in medical imaging, such as a dearth of high-quality picture datasets and labeled data as compared to other high-quality natural image datasets.
+Adversarial defenses might not have a limited impact.
+The results of the study highlight the need for greater care in designing DNNs for medical imaging and their practical applications.
+Figure 8: A demonstration of how adversarial attacks against different medical AI systems could be carried out without requiring any openly dishonest data manipulation [53].